OpenStudy (anonymous):

if x≠0, then u/x+ 5u/x-u/5x= A. 7x/5u B. 5u/7x C. 29u/5x D. 31u/5x

OpenStudy (tkhunny):

It's fractions. Add them? Get the denominators the same and add the numerators.

OpenStudy (anonymous):

so is 7 x for the denominators? what about the u? they look too spreadeed out

OpenStudy (samanthagreer):

i belive it is c

OpenStudy (anonymous):

why is it c?

OpenStudy (anonymous):

how did it became like 29 u if there is only like 9 u

OpenStudy (tkhunny):

Is it this? \(\dfrac{u}{x}+ \dfrac{5u}{x}-\dfrac{u}{5x}\) Find a common denominator. We're just adding fractions.

OpenStudy (anonymous):

So if we are just adding it wouldnt be 7u/7x?

OpenStudy (anonymous):

Okay everything here is good :)

OpenStudy (tkhunny):

?? \(\dfrac{u}{x} + \dfrac{5u}{x} - \dfrac{u}{5x}\) Common Denominator \(\dfrac{5u}{5x} + \dfrac{25u}{5x} - \dfrac{u}{5x}\) Add \(\dfrac{5u + 25u - u}{5x} = \dfrac{29u}{5x}\) Now, how are you feeling about anything that uses the number 7?

OpenStudy (anonymous):

where did the 25 appear from?

OpenStudy (carlyleukhardt):

Its c.

OpenStudy (anonymous):

I still dont understand

OpenStudy (welshfella):

the lowest common multiple of x,x and 5x is 5x so you can do this by making the denominators in the 3 fractions the same for the first fraction u/x you multiply the u and the x by 5 giving 5u/5x

OpenStudy (welshfella):

do you follow that?

OpenStudy (lochana):

@Kfsv10 have you found a solution?

OpenStudy (anonymous):

So it could become in any number?

OpenStudy (lochana):

yes. It is just solving fractions. for example x(1/3 + 2/3) is same as x/3 + 2x/3. so u/x+ 5u/x-u/5x = u/x(1/1 + 5/1 - 1/5) = u/x(29/5) which is 29u/5x

OpenStudy (welshfella):

you need to convert 5u/x to a fraction with 5x as the denominator in the same way as above . The last fraction is u/5x has a denominator of 5x already.